# Solver-Free Decision-Focused Learning for Linear Optimization Problems

### v1

@@ 1 Introduction
Traditionally, the machine learning model is trained to maximize accuracy (e.g., by minimizing the mean squared error over all its predictions). However, this may lead to suboptimal decisions. This is because, while perfect predictions lead to perfect decisions, different kinds of imperfect predictions affect downstream decision-making in different ways leaving_the_nest . Decision-focused learning (DFL) addresses this by training models specifically to make predictions that lead to good decisions. As numerous works have shown, this generally improves decision quality (optnet, ; donti2017task, ; spo, ; ltr, ; NCE, ; blackbox, ; melding, ).
@@ 2.1 Predict-then-optimize
with $c\in\mathbb{R}^{n}$, $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $m\leq n$. We denote the set of optimal solutions as $Z^{\star}(c)\subseteq\mathbb{R}^{n}$ is the set of optimal solutions, and a particular optimal solution as $z^{\star}(c)\in Z^{\star}(c)$. The parametric LP is written in standard form (i.e., with equality constraints), without loss of generality: if an LP is not initially in standard form, it can be converted by introducing slack variables. We further assume that the rows of the constraint matrix $A$ are linearly independent. If not, the problem includes redundant (implied) constraints that can be removed without affecting the feasible region. Finally, we assume that the feasible region $\{z\in\mathbb{R}^{n}\mid Az=b,\,z\geq 0\}$ is non-empty and bounded.
@@
Unlike conventional regression, the objective in training is not to maximize the accuracy of predicted costs $\hat{c}$. Rather, the aim is to make predictions that maximize the quality of the resulting decisions. This is measured by the regret, which expresses the suboptimality of the made decisions $z^{\star}(\hat{c})$ with respect to true costs $c$ (lower is better).
@@ 2.2 Linear programming preliminaries
Given a matrix $A\in\mathbb{R}^{m\times n}$ with full row rank, a basis is a set of $m$ linearly independent columns of $A$. These columns form an invertible matrix $B\in\mathbb{R}^{m\times m}$. The variables corresponding to the columns in the basis are called basic variables; the remaining $n-m$ variables are called non-basic variables.
@@ 3 Related work
A lot of work on DFL has focused specifically on LPs and combinatorial problems, as handling these problems is particularly challenging. This is because they make the gradient of any loss that depends on $z^{\star}(\hat{c})$ (like the regret) zero almost everywhere. This can be seen when applying the chain rule:
@@
The second factor $\frac{\partial z^{\star}(\hat{c})}{\partial\hat{c}}$ expresses the change in $z^{\star}(\hat{c})$ when $\hat{c}$ changes infinitesimally. However, for LPs and combinatorial problems, this factor is zero almost everywhere, and nonexistent otherwise. In other words, a small change in the problem’s parameters $\hat{c}$ either does not change its solution $z^{\star}(\hat{c})$, or changes it discontinuously (leading to nonexistent gradients). Most work on DFL has focused on circumventing this obstacle. Three general types of approaches can be distinguished. We briefly discuss them here, but refer the reader to dfl_survey for a comprehensive overview of the field.
@@ 4 Solver-free training by contrasting adjacent vertices
Figure 1: Optimal solution $z^{\star}$ has two adjacent vertices, $z_{adj}^{(1)}$ and $z_{adj}^{(2)}$. Geometrically, these vertices define the facets of the optimality cone of $z^{\star}$. All cost vectors $c$ for which $-c$ lies within this cone lead to $z^{\star}$. Cost vector $\hat{c}$ is predicted. $\hat{c}$ correctly prefers $z^{\star}$ to $z_{adj}^{(1)}$ (i.e., $\hat{c}^{\top}z^{\star}<\hat{c}^{\top}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}z_{adj}^{(1)}}$), and thus lies on the correct side of the corresponding facet. This adds no penality to $\mathcal{L}_{\text{AVA}}$. However, $\hat{c}$ wrongly prefers $z_{adj}^{(2)}$ to $z^{\star}$, and thus lies on the wrong side of the corresponding facet. This adds $\hat{c}^{\top}z^{\star}-\hat{c}^{\top}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{adj}^{(2)}}$ to $\mathcal{L}_{\text{AVA}}$.
@@ 4.1 Loss based on adjacent vertices
We construct a loss function based on the following proposition (which we prove in the appendix).
@@
The geometric interpretation of the lava loss is shown in Figure 1. The loss evaluates how well the predicted cost vector $\hat{c}$ distinguishes a true optimal solution $z^{\star}(c)$ from its adjacent vertices. Specifically, it penalizes cases where an adjacent solution $z_{adj}\in Z_{adj}$ appears more favorable than the optimal one under $\hat{c}$. For each adjacent vertex, the loss increases proportionally to the margin by which $c^{\top}z^{\star}(c)$ is worse than $c^{\top}z_{adj}-\epsilon$. Once $z^{\star}(c)$ is better than $z_{adj}$ by at least $\epsilon$, no further gain is enforced, reflecting the fact that, in Proposition 4.1, the precise margin by which $c^{\top}z\leq c^{\top}z_{adj}$ is irrelevant. In case of multiple optimal solutions (i.e., $|Z^{\star}(c)|>1$), the loss is more conservative, and is minimized only for a proper subset of cost vectors that leads to $z^{\star}(c)$ as optimal solution.
The value of $\epsilon$ defines a margin for how much better the optimal solution must be before further improvements are not rewarded. In principle, $\epsilon=0$ is most in line with the optimality condition expressed in Proposition 4.1. However, in practice, using a small positive value (e.g., $\epsilon=0.1$, which we use in our experiments) introduces a buffer zone that tends to lead to smoother and more stable learning, likely by reducing sensitivity to minor fluctuations near the decision boundary.
@@
    It is differentiable with respect to $\hat{c}$ and offers non-zero gradients whenever $\hat{c}$ does not yet lead to the ground-truth optimal solution $z^{\star}(c)$.
@@
    It is convex, ensuring that any local minimum is also a global minimum. To see why, note that both terms in $\hat{c}^{\top}z^{\star}(c)-\hat{c}^{\top}z_{adj}$ are affine functions of $\hat{c}$, and thus their difference is also affine (and hence convex). Taking the maximum of this affine function and constant $-\epsilon$ leads to a hinge-like function that remains convex. And since the sum of convex functions remains convex, the overall loss remains convex when taking the sum over all $z_{adj}\in Z_{adj}(z^{\star}(c))$.
@@
    It is computationally efficient to evaluate, as it does not require computing $z^{\star}(\hat{c})$ through a call to a solver. Instead, it only involves dot products between $\hat{c}$ and a set of precomputed solution vectors, making it highly efficient.
The lava loss is somewhat related to the recently proposed CaVE tang2024cave , particularly in terms of underlying objectives. The CaVE loss is the cosine distance between $\hat{c}$ and its projection onto the optimality cone of solution $z^{\star}(c)$. Computing this projection involves solving a QP. Instead, lava penalizes the ‘violation’ of each facet of the cone separately, as illustrated in Figure 1. This requires no solve calls.
lava is also related to the NCE loss from contrastive , but whereas NCE requires an arbitrary set of feasible solutions constructed via solve calls on predicted cost vectors during training, lava contrasts the optimal solution with its adjacent vertices, which can be precomputed without solving. Additionally, lava is a hinge-like loss thresholded at $-\epsilon$, aligning with the LP optimality condition (Proposition 4.1), whereas NCE seeks to increase the difference in objective values as much as possible. This thresholding is crucial to lava’s performance, as demonstrated in an ablation in the appendix.
@@ 4.2 Precomputing the adjacent vertices
1:Input: Basic feasible solution $z^{\star}$, basic indices $B(1),\ldots,B(m)$, non-basic indices $N(1),\ldots,N(n-m)$, constraint matrix $A$
@@
[…]
8:Dequeue a basis $B_{curr}$ and corresponding $N_{curr}$ from $Queue$
9:Compute directions of movement $D=-B_{curr}^{-1}N_{curr}$Foreach non-basic variable $z^{\star}_{j}$ as entering variable
10:$d^{\prime}=D_{\ast j}$ is the direction of movement of the basic variables when increasing $z^{\star}_{j}$ by $1$
[…]
13:Set direction $d_{B_{curr}(k)}=d^{\prime}_{k}$ for basic variable $B_{curr}(k)$EndFor
[…]
15:Identify $B_{curr}(i)$ as the leaving basic variable according to the TNP rule
16:Construct new basis $B_{new}$ and corresponding $N_{new}$ by replacing $B(i)$ with $N(j)$If$(B_{new},N_{new})\notin Visited$
[…]
We now outline the procedure that precomputes the adjacent vertices $Z_{adj}$ of a vertex $z^{\star}$ in Algorithm 1. The algorithm takes as input a basic feasible solution $z^{\star}$, the constraint matrix $A$, and a basis of $z^{\star}$ (expressed through the indices of the basic and non-basic variables). It outputs the set $Z_{adj}$ of all adjacent vertices. The algorithm starts by constructing the initial basis matrix $B$ (and the associated matrix of non-basic columns $N$), and uses these to perform pivots, as performed in the (revised) simplex method bertsimas1997introduction ; dantzig1951maximization . Here, a pivot refers to the swapping of a basic variable and a non-basic variable. The progression through the algorithm differs depending on whether $z^{\star}$ is a nondegenerate or degenerate solution. We now discuss each case separately.
@@ 4.2.1 Base case: nondegenerate vertices
When $z^{\star}$ is nondegenerate, it has a unique associated basis, and every adjacent vertex can be obtained through a single pivot operation (i.e., by replacing one basic variable with a non-basic one). To do so, $D=-B^{-1}N$ is computed at line 9. Each column $D_{\ast j}$ expresses how the basic variables must change to maintain $Az=b$ when non-basic variable $z_{j}$ is increased by $1$. The algorithm then loops through the non-basic variables, and considers each $z^{\star}_{j}$ in turn as the variable that enters the basis (line 9). This differs from the simplex method, which selects a single entering variable heuristically.
For each entering variable, the algorithm must determine the corresponding variable that should leave the basis. To do so, it computes the maximum step size $\theta^{\star}$ that can be taken in direction $d^{\prime}=D_{\ast j}$ without violating the nonnegativity constraints $z\geq 0$. This is computed using the standard minimum ratio test $\theta^{\star}=\min_{i:d^{\prime}_{i}<0}\left(-\nicefrac{{z^{\star}_{i}}}{{d^{\prime}_{i}}}\right)$ bertsimas1997introduction ; dantzig1951maximization . Since $z^{\star}$ is nondegenerate, $\theta^{\star}>0$ will hold, and the if-block at line 11 will be entered. The corresponding adjacent vertex is now given by $z^{\star}+\theta^{\star}d$, where $d\in\mathbb{R}^{n}$ is obtained by setting the entry corresponding to the entering variable $j$ to $1$, and setting the value of each basic variable index $B(i)$ to $d^{\prime}_{i}$, with all other entries remaining zero (lines 12-14).
Because of nondegeneracy, the else-block at line 14 is never entered, and all adjacent vertices are found after a single iteration of the while-loop.
@@ 4.2.2 Edge case: degenerate vertices
In the degenerate case, multiple bases correspond to $z^{\star}$. This makes efficiently computing all adjacent vertices much more challenging, as they cannot all be obtained by performing pivots on any single basis $B$ of $z^{\star}$. In this case, some of the basis pivots performed in Algorithm 1 have a step size of $\theta^{\star}=0$, and result in the same vertex $z$. These pivots thus produce other bases associated with $z$, which subsequently have to be explored. Thus, to ensure that all adjacent vertices of $z$ are identified, one in principle needs to consider all pivot operations across all bases associated with $z$. However, as shown in upper_bound_reference and degeneracy_graphs , this quickly becomes intractable.
Some work has aimed to address this, primarily between the late 1970s and early 1990s, but has since remained relatively obscure, receiving limited attention in recent research. For an overview, we refer the reader to degeneracy_tutorial ; degeneracy_updated_survey ; solved_and_open_degeneracy_problems . Particularly relevant to this paper is the work by Geue TNP-N-Tree and Kruse degeneracy_graphs , which showed that to compute all adjacent vertices of a degenerate solution, not all bases of that solution need to be explored. It suffices to only consider those bases that can be reached through lexicographic pivoting lexicographic_pivoting_rule . To further reduce the required number of pivots, Geue later introduced a dedicated Transition Node Pivoting (TNP) rule TNP , which we employ in Algorithm 1. More information on the TNP rule can be found in the appendix.
To systematically explore bases using the TNP rule, we use the Queue and Visited data structures initialized in lines 6 and 7, respectively. When an entering variable results in a step size $\theta^{\star}=0$ in line 11, the algorithm applies the TNP rule to select a leaving variable (line 15). If the resulting basis is not in Visited yet, it is added to the Queue for later exploration. The Visited set ensures that each basis is processed only once, preventing redundant pivot operations.

### v2

@@ 1 Introduction
Traditionally, the machine learning model is trained to maximize accuracy (e.g., by minimizing the mean squared error over all its predictions). However, this may lead to suboptimal decisions. This is because, while perfect predictions lead to perfect decisions, different kinds of imperfect predictions affect downstream decision-making in different ways. Decision-focused learning (DFL) addresses this by training models specifically to make predictions that lead to good decisions. As numerous works have shown, this generally improves decision quality (optnet, ; donti2017task, ; spo, ; ltr, ; NCE, ; blackbox, ; melding, ).
@@ 2.1 Predict-then-optimize
with $c\in\mathbb{R}^{n}$, $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $m\leq n$. We denote the set of optimal solutions as $Z^{\star}(c)\subseteq\mathbb{R}^{n}$, and a particular optimal solution as $z^{\star}(c)\in Z^{\star}(c)$. The parametric LP is written in standard form (i.e., with equality constraints), without loss of generality: if an LP is not initially in standard form, it can be converted by introducing slack variables. We further assume that the rows of the constraint matrix $A$ are linearly independent. If not, the problem includes redundant (implied) constraints that can be removed without affecting the feasible region. Finally, we assume that the feasible region $\{z\in\mathbb{R}^{n}\mid Az=b,\,z\geq 0\}$ is non-empty and bounded.
@@
[…]
@@ 2.2 Linear programming preliminaries
Given a matrix $A\in\mathbb{R}^{m\times n}$ with full row rank, a basis is a set of $m$ linearly independent columns of $A$. These columns form an invertible matrix $B\in\mathbb{R}^{m\times m}$. The variables corresponding to the columns in the basis are called basic variables; the remaining nmn-m variables are called non-basic variables.
@@ 3 Related work
[…]
@@
[…]
@@ 4 Solver-free training by contrasting adjacent vertices
Figure 1: Optimal solution $z^{\star}$ has two adjacent vertices, $z_{adj}^{(1)}$ and $z_{adj}^{(2)}$. Geometrically, these vertices define the facets of the optimality cone of $z^{\star}$. All cost vectors $c$ for which c-c lies within this cone lead to $z^{\star}$. Cost vector $\hat{c}$ is predicted. $\hat{c}$ correctly prefers $z^{\star}$ to $z_{adj}^{(1)}$ (i.e., $\hat{c}^{\top}z^{\star}<\hat{c}^{\top}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}z_{adj}^{(1)}}$), and thus lies on the correct side of the corresponding facet. This adds no penality to $\mathcal{L}_{\text{AVA}}$. However, $\hat{c}$ wrongly prefers $z_{adj}^{(2)}$ to $z^{\star}$, and thus lies on the wrong side of the corresponding facet. This adds $\hat{c}^{\top}z^{\star}-\hat{c}^{\top}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{adj}^{(2)}}$ to $\mathcal{L}_{\text{AVA}}$.
@@ 4.1 Loss based on adjacent vertices
We construct a loss function based on the following proposition, which follows from the convexity of the feasible polytope (as proven in Appendix A).
@@
The geometric interpretation of the lava loss is shown in Figure 1. The loss evaluates how well the predicted cost vector $\hat{c}$ distinguishes a true optimal solution $z^{\star}(c)$ from its adjacent vertices. Specifically, it penalizes cases where an adjacent solution $z_{adj}\in Z_{adj}$ appears more favorable than the optimal one under $\hat{c}$. Geometrically, this occurs when the predicted cost vector $\hat{c}$ falls outside the optimality cone of $z^{\star}(c)$, defined as the set of cost vectors for which $z^{\star}(c)$ remains optimal. For each adjacent vertex, the loss increases proportionally to the margin by which $c^{\top}z^{\star}(c)$ is worse than $c^{\top}z_{adj}-\epsilon$. Once $z^{\star}(c)$ is better than $z_{adj}$ by at least $\epsilon$, no further gain is enforced, reflecting the fact that, in Proposition 4.1, the precise margin by which $c^{\top}z\leq c^{\top}z_{adj}$ is irrelevant. In case there are multiple optimal solutions for ground-truth $c$ (i.e., $|Z^{\star}(c)|>1$), the loss is more conservative, and is minimized only for a proper subset of cost vectors that leads to $z^{\star}(c)$ as optimal solution.
The value of $\epsilon$ defines a margin for how much better the optimal solution must be before further improvements are not rewarded. In principle, $\epsilon=0$ is most in line with the optimality condition expressed in Proposition 4.1. However, after meeting the optimality condition, subsequent updates during training (e.g., from other instances) may shift the predicted cost vector slightly such that an adjacent vertex becomes preferred again. To prevent this, using a small positive value (e.g., $\epsilon=0.1$, which we use in our experiments) introduces a buffer zone that tends to lead to smoother and more stable learning, likely by reducing sensitivity to minor fluctuations near the decision boundary. In Appendix B, we provide a visualization of how this margin parameter affects the learning curve.
@@
    It is differentiable with respect to $\hat{c}$ almost everywhere and offers non-zero gradients whenever $\hat{c}$ does not yet lead to the ground-truth optimal solution $z^{\star}(c)$.
@@
    It is convex, ensuring that any local minimum is also a global minimum. To see why, note that both terms in $\hat{c}^{\top}z^{\star}(c)-\hat{c}^{\top}z_{adj}$ are affine functions of $\hat{c}$, and thus their difference is also affine (and hence convex). Taking the maximum of this affine function and constant $\epsilon$ leads to a hinge-like function that remains convex. And since the sum of convex functions remains convex, the overall loss remains convex when taking the sum over all $z_{adj}\in Z_{adj}(z^{\star}(c))$.
@@
[…]
The lava loss is somewhat related to the recently proposed CaVE tang2024cave , particularly in terms of underlying objectives. The CaVE loss is the cosine distance between $\hat{c}$ and its projection onto the optimality cone of solution $z^{\star}(c)$. Computing the projection involves solving a QP. Instead, lava penalizes the “violation” of each facet of the cone separately, as illustrated in Figure 1. This requires no solve calls. Another important difference is that CaVE represents the optimality cone through its extreme rays, which correspond to the binding constraints in $z^{\star}(c)$. In contrast, lava operates on the cone’s facets, each defined by the hyperplane separating $z^{\star}(c)$ from an adjacent vertex $z_{adj}$.
lava is also related to the NCE loss from NCE , but whereas NCE requires an arbitrary set of feasible solutions constructed via solve calls on predicted cost vectors during training, lava contrasts the optimal solution with its adjacent vertices, which can be precomputed without solving. Additionally, lava is a hinge-like loss thresholded at $\epsilon$, aligning with the LP optimality condition (Proposition 4.1), whereas NCE seeks to increase the difference in objective values as much as possible. This thresholding is crucial to lava’s performance, as demonstrated in an ablation in Appendix B.
Note that lava is not differentiable everywhere, because each of its $\max(\hat{c}^{\top}z^{\star}(c)-\hat{c}^{\top}z_{adj},-\epsilon)$ terms has a hinge-like form, which is not smooth at the kink where $\hat{c}^{\top}z^{\star}(c)-\hat{c}^{\top}z_{adj}=-\epsilon$. However, these kink points constitute a measure-zero set, meaning that the probability of the predicted cost vector $\hat{c}$ ending up exactly at a kink point is zero. Consequently, lava can be optimized reliably using standard gradient-based methods.
Also note that, although the zero vector technically lies within every optimality cone and could in principle minimize lava when the margin $\epsilon$ is zero, this does not occur in practice. Standard random initialization of the predictive model ensures that the predicted cost vector starts away from zero, and gradient updates induced by lava push predictions toward differences between optimal and adjacent solutions, rather than toward the origin. While pathological scenarios could be constructed in which all gradients align exactly toward zero, such cases are extremely unlikely and have not been observed in practice.
We presented lava in the context of LPs. However, it also extends to integer linear programs (ILPs) and mixed-integer linear programs (MILPs) by operating on their LP relaxations. The effect of doing so depends on the properties of the original problem. For binary ILPs, each optimality cone in the LP relaxation is contained within the corresponding cone of the binary ILP tang2024cave . This only makes lava somewhat conservative: it will always push predicted cost vectors toward the correct optimality cone, but it may push them further into the cone than strictly needed. For non-binary ILPs and MILPs, this containment does not necessarily hold, and the performance of using lava largely depends on the integrality gap of the relaxation.
@@ 4.2 Precomputing the adjacent vertices
1: Input: Basic feasible solution $z^{\star}$, basic indices $B(1),\ldots,B(m)$, non-basic indices $N(1),\ldots,N(n-m)$, constraint matrix AA
@@
7: $Visited\leftarrow\{(B,N)\}$
8: while QueueQueue is not empty do
9:   Dequeue a basis $B_{curr}$ and corresponding $N_{curr}$ from QueueQueue
10:   Compute directions of movement $D=-B_{curr}^{-1}N_{curr}$
11:   for each non-basic variable $z^{\star}_{j}$ as entering variable do
12:    $d^{\prime}=D_{\ast j}$ is the direction of movement of the basic variables when increasing $z^{\star}_{j}$ by 11
13:    Perform minimum ratio test $\theta^{\star}=\min\limits_{\{i\mid d_{i}<0\}}\left(-\frac{z^{\star}_{B_{curr}(i)}}{d_{i}}\right)$
14:    if $\theta^{\star}>0$ then $\triangleright$ Nondegenerate pivot
15:      Let $d$ be the unit vector $e_{j}\in\mathbb{R}^{n}$
16:      for $k=1\ldots m$ do
17:       Set direction $d_{B_{curr}(k)}=d^{\prime}_{k}$ for basic variable $B_{curr}(k)$
18:      Add new adjacent vertex $Z_{adj}=Z_{adj}\cup\{z^{\star}+\theta^{\star}d\}$
19:    else $\triangleright$ Degenerate pivot
20:      Identify $B_{curr}(i)$ as the leaving basic variable according to the TNP rule
21:      Construct new basis $B_{new}$ and corresponding $N_{new}$ by replacing $B(i)$ with $N(j)$
22:      if $(B_{new},N_{new})\notin Visited$ then
23:       Add $(B_{new},N_{new})$ to QueueQueue and to VisitedVisited
24: return $Z_{adj}$
We outline the procedure that precomputes the adjacent vertices $Z_{adj}$ of a vertex $z^{\star}$ in Algorithm 1. The algorithm takes as input a basic feasible solution $z^{\star}$, the constraint matrix $A$, and a basis of $z^{\star}$ (expressed through the indices of the basic and non-basic variables). It outputs the set $Z_{adj}$ of all adjacent vertices. The algorithm starts by constructing the initial basis matrix $B$ (and the associated matrix of non-basic columns $N$), and uses these to perform pivots, as performed in the (revised) simplex method bertsimas1997introduction ; dantzig1951maximization . Here, a pivot refers to the swapping of a basic variable and a non-basic variable. The progression through the algorithm differs depending on whether $z^{\star}$ is a nondegenerate or degenerate solution. We now discuss each case separately.
@@ 4.2.1 Base case: nondegenerate vertices
When $z^{\star}$ is nondegenerate, it has a unique associated basis, and every adjacent vertex can be obtained through a single pivot operation (i.e., by replacing one basic variable with a non-basic one). To do so, $D=-B^{-1}N$ is computed at line 10 of Algorithm 1. Each column $D_{\ast j}$ expresses how the basic variables must change to maintain $Az=b$ when non-basic variable $z_{j}$ is increased by $1$. The algorithm then loops through the non-basic variables, and considers each $z^{\star}_{j}$ in turn as the variable that enters the basis (line 11). This differs from the simplex method, which selects a single entering variable heuristically.
For each entering variable, the algorithm must determine the corresponding variable that should leave the basis. To do so, it computes the maximum step size $\theta^{\star}$ that can be taken in direction $d^{\prime}=D_{\ast j}$ without violating the nonnegativity constraints $z\geq 0$. This is computed using the standard minimum ratio test $\theta^{\star}=\min_{i:d^{\prime}_{i}<0}\left(-\nicefrac{{z^{\star}_{i}}}{{d^{\prime}_{i}}}\right)$ bertsimas1997introduction ; dantzig1951maximization . Since $z^{\star}$ is nondegenerate, $\theta^{\star}>0$ will hold, and the if-block at line 14 will be entered. The corresponding adjacent vertex is now given by $z^{\star}+\theta^{\star}d$, where $d\in\mathbb{R}^{n}$ is obtained by setting the entry corresponding to the entering variable $j$ to $1$, and setting the value of each basic variable index $B(i)$ to $d^{\prime}_{i}$, with all other entries remaining zero (lines 15-18).
Because of nondegeneracy, the else-block at line 19 is never entered, and all adjacent vertices are found after a single iteration of the while-loop.
@@ 4.2.2 Edge case: degenerate vertices
In the degenerate case, multiple bases correspond to $z^{\star}$. This makes efficiently computing all adjacent vertices much more challenging, as they cannot all be obtained by performing pivots on any single basis $B$ of $z^{\star}$. In this case, some of the basis pivots performed in Algorithm 1 have a step size of $\theta^{\star}=0$, and result in the same vertex $z^{\star}$. These pivots thus produce other bases associated with $z^{\star}$, which subsequently have to be explored. Thus, to ensure that all adjacent vertices of $z^{\star}$ are identified, one in principle needs to consider all pivot operations across all bases associated with $z^{\star}$. However, as shown in upper_bound_reference and degeneracy_graphs , this quickly becomes intractable.
Some work has aimed to address this, primarily between the late 1970s and early 1990s, but has since remained relatively obscure, receiving limited attention in recent research. For an overview, we refer the reader to degeneracy_tutorial ; degeneracy_updated_survey ; solved_and_open_degeneracy_problems . Particularly relevant to this paper is the work by Geue TNP-N-Tree and Kruse degeneracy_graphs , which showed that to compute all adjacent vertices of a degenerate solution, not all bases of that solution need to be explored. It suffices to only consider those bases that can be reached through lexicographic pivoting lexicographic_pivoting_rule . To further reduce the required number of pivots, Geue later introduced a dedicated Transition Node Pivoting (TNP) rule TNP , which we employ in Algorithm 1. More information on the TNP rule can be found in Appendix C.
To systematically explore bases using the TNP rule, we use the Queue and Visited data structures initialized in lines 6 and 7, respectively. When an entering variable results in a step size $\theta^{\star}=0$ in line 13, the algorithm applies the TNP rule to select a leaving variable (line 20). If the resulting basis is not in Visited yet, it is added to the Queue for later exploration. The Visited set ensures that each basis is processed only once, preventing redundant pivot operations.